Ethiopia has a GDP of $8 billion (measured in U.S. dollars) and a population of 55 million. Costa Rica has a GDP of $9 billion (measured in U.S. dollars) and a population of 4 million. Calculate per person GDP for each country.

Group of answer choices

Ethiopia = $14.50 Costa Rica = $2250.00

Ethiopia = $14.50 Costa Rica = $225.00

Ethiopia = $145.00 Costa Rica = $2250.00

Ethiopia = $1450.00 Costa Rica = $22,500.00
Answer #1

Answer

Option 3

Ethiopia = $145.00 Costa Rica = $2250.00

==============

Per capita GDP=GDP/Population

Working in million

Per capita GDP of Ethiopia =8000/55=145.454545=145.50

Per capita GDP of Costa Rica =9000/4=2250=2250.00